40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 4 10:30:37 2016, ericq, Update, General, Preliminary vent plans 
    Reply  Tue Jul 5 17:14:02 2016, ericq, Update, General, Vent progress noOsemSide.JPGOSEMside.JPG
       Reply  Tue Jul 5 19:10:41 2016, rana, Update, General, Vent progress jumpEvent20151105.png
          Reply  Tue Jul 5 21:51:35 2016, ericq, Update, General, Vent progress 
             Reply  Tue Jul 5 23:38:42 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation image.jpeg
                Reply  Wed Jul 6 22:58:01 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                   Reply  Thu Jul 7 00:25:07 2016, ericq, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                      Reply  Thu Jul 7 10:49:03 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation UL_glitchMon_Striptool.png
                         Reply  Fri Jul 8 11:35:45 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation coilGlitchMon.png
                            Reply  Fri Jul 8 11:48:09 2016, Koji, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                               Reply  Fri Jul 8 14:03:26 2016, ericq, Update, General, ETMX removed from suspension zoom.jpg
                                  Reply  Fri Jul 8 16:33:26 2016, ericq, Update, General, ETMX removed from suspension 
                                     Reply  Fri Jul 8 20:44:09 2016, Koji, Update, General, ETMX removed from suspension 6x
                                        Reply  Sun Jul 10 15:38:01 2016, ericq, Update, General, ETMX in guiderod/standoff gluing fixture 6x
                                           Reply  Tue Jul 12 16:08:00 2016, ericq, Update, General, ETMX ruby guided gluing 
                                              Reply  Wed Jul 13 21:13:51 2016, Koji, Update, General, ETMX guide rod gluing / ETMY Magnet gluing 13x
                                                 Reply  Thu Jul 14 22:29:57 2016, ericq , Update, General, ETMX guide rod gluing / ETMY Magnet gluing 
                                                    Reply  Fri Jul 15 16:23:51 2016, ericq , Update, General, Gluing setbacks IMG_2768.JPGIMG_2771.JPG
                                                       Reply  Tue Jul 19 16:39:29 2016, Lydia, Update, General, ETMX magnet gluing/guiderod excess glue removed P7190201.JPGP7190203.JPG
                                                          Reply  Wed Jul 20 13:58:55 2016, Steve, Update, SUS, SUS damping out of vac chamber ETMXrack.jpgfromETMX-satBox.jpgsusDampingSatboxCab.jpgsusDampingSatboxCabl.jpg
                                                          Reply  Thu Jul 21 17:01:59 2016, ericq, Update, General, ETMX magnet gluing/guiderod excess glue removed 
                                                             Reply  Fri Jul 22 23:03:23 2016, gautam, Update, General, ETMX suspension - progress IMG_2828.JPGIMG_2826.JPGIMG_2821.JPGIMG_2822.JPGIMG_2823.JPG
                                                                Reply  Sun Jul 24 03:43:56 2016, ericq, Update, General, Magnet positioning 
                                                                Reply  Mon Jul 25 13:01:49 2016, ericq, Update, General, ETMX suspension - progress IMG_2829.JPGIMG_2830.JPGIMG_2831.JPGIMG_2832.JPGIMG_2833.JPG
                                                                   Reply  Mon Jul 25 13:47:19 2016, rana, Update, General, ETMX suspension - progress 
                                                                   Reply  Mon Jul 25 17:01:54 2016, Steve, Update, SUS, ETMX suspension - progress ITMYside2011.JPG
                                                                   Reply  Mon Jul 25 20:37:46 2016, gautam, Update, General, ETMX suspension - Side OSEM test osem_test_2016_07_25.png
                                                                      Reply  Tue Jul 26 18:07:57 2016, gautam, Update, General, ETMX suspension - Further OSEM Tests OSEM_spectra.pdf
                                                                         Reply  Thu Jul 28 03:22:28 2016, ericq, Update, General, ETMX magnets re-reglued 
                                                                            Reply  Thu Jul 28 13:49:44 2016, gautam, Update, General, ETMX magnets re-reglued ETMX_sideMagnet.JPGETMX_sideMagnet_zoomed.JPG
                                                                               Reply  Thu Jul 28 17:48:17 2016, gautam, Update, General, ETMY LR magnet broke off 
                                                                                  Reply  Thu Jul 28 22:43:03 2016, rana, Update, General, ETMY UL magnet broke off 
                                                                                  Reply  Wed Aug 3 10:52:03 2016, Steve, Update, General, ETMY soaking dish SD1.jpgSD2.jpgSD3.jpg
                                                                               Reply  Fri Jul 29 03:32:30 2016, ericq, Update, General, ETMX suspended once more 
                                                                                  Reply  Fri Jul 29 13:17:34 2016, Koji, Update, General, Oven 
                                              Reply  Thu Jul 14 01:00:33 2016, Koji, Update, General, ETMX ruby guided gluing ruby.jpg
                                  Reply  Thu Jul 28 16:43:01 2016, Lydia, Update, General, ETMX aluminium standoff groove condition wire_damage_4.jpgwire_damage_zoom.jpgno_wire.jpgcomparison_2.jpg
                            Reply  Fri Jul 8 21:22:38 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
             Reply  Wed Jul 6 09:34:26 2016, Steve, Update, General, PRM sensing voltage LR PRM_LR.png
             Reply  Wed Jul 6 19:36:45 2016, Koji, Update, General, SUS Vmon HPF.png160706_ITMX_VMON2.pdf160706_ITMX_VMON1.pdf
                Reply  Thu Jul 7 15:23:39 2016, ericq, Update, General, SUS Vmon coil_drive.png
                   Reply  Thu Jul 7 16:05:55 2016, Koji, Update, General, SUS Vmon 
       Reply  Thu Jul 7 17:12:50 2016, Johannes, Update, General, Vent progress IMG_2621.JPGIMG_2623.JPGIMG_2628.JPGIMG_2595.JPG
          Reply  Tue Jul 12 23:51:16 2016, Johannes, Update, General, Vent progress - ETMY inspection IMG_2679.JPGIMG_2673.JPG
             Reply  Wed Jul 13 00:38:25 2016, Johannes, Update, General, Vent progress - ETMY attempted repositioning 
                Reply  Tue Jul 19 13:21:42 2016, Johannes, Update, General, Vent progress - ETMY attempted repositioning 
       Reply  Sun Jul 10 17:33:00 2016, ericq, Update, General, Vent progress 
          Reply  Mon Jul 11 15:13:22 2016, gautam, Update, General, Vent progress: in-date First Contact procured image.jpeg
Message ID: 12340     Entry time: Tue Jul 26 18:07:57 2016     In reply to: 12335     Reply to this: 12345
Author: gautam 
Type: Update 
Category: General 
Subject: ETMX suspension - Further OSEM Tests 

Today, we did the following:

  • Once again, inserted all four face OSEMs till the sensor voltage readouts were approximately half their saturation value. The presence of some ferromagnetic material in the Honeywell components makes this tricky as each coil is coupled to the other three, but we were able to converge to a point where all the voltage readouts were oscillating around a mean value of ~40-60% of their maximum value, with all the damping loops OFF.
  • Turned on all damping loops, and verified that the OSEM positioning was indeed such that the sensor readout is nominally around 50% of the saturation value. The air buffeting around the clean bench means that the damping isn't nearly as effective as it is inside the vacuum chamber.
  • Attempted to increase the gain on the damping loops - we first switched out the Chebyshev low-pass filter in all the damping loops for something a little less aggressive, to allow us to turn up the gain. However, this experiment wasn't a success, when we turned the damping loops on, they were ringing the optic up.
  • At this point, Eric checked the offset sliders (summed in via the slow system) and saw that they were not zero. We zeroed these, but naturally, they destroyed the OSEM positioning equilibrium we had established earlier. So we had to go back and re-position the OSEMs
  • After re-centering the face OSEM magnets relative to the LED-PD pair, we insertd the side OSEM such that the side magnet completely occluded the PD. Interestingly, Eric noticed that the magnetic attraction between OSEM and magnets conspired to center the side magnet fairly well in the side OSEM, when it completely blocked the PD. However, when he returned the side OSEM coil to its nominal operating position of approximately half-blocking the PD, some minor misalignment was re-introduced (i.e. even when the optic was swinging mostly along the axis of the side OSEM, the voltage readout did not quite go down to 0). 
  • We then decided to compare the spectra for the error signals for the 4 DOFs with the current configuration (i.e. suspension clamped down to table top, optic freely hanging, all OSEMs reasonably well centered, and with the ETMX SUS model reverted to its normal state) to some reference (see Attachment #1). I initially thought I would wait for the optic to settle down a bit more before taking the spectra, but it doesn't seem to be showing any signs of getting any quieter in the last one hour. In Attachment #1, I have plotted as reference the spectra of the error signals from the early hours of 4 July 2016, at which point we were at atmosphere but the heavy doors were not yet off, so this is not really a fair comparison, but we don't really have a period in which the optic was exposed to the atmosphere and with the OSEMs in place, at least from this vent. Colors are identical for a given DOF, with todays trace as a solid line, and the reference dashed.
  • We did not check the room available to install some shimming piece of metal in the side OSEM holder, as a possible solution to solve the misalignment problem. Steve has already found pieces of varying thickness, and they are soaking in acetone right now, we plan to air bake them tomorrow. 

I will have another look at the spectra tomorrow morning, to see if the damping improves overnight. 

Attachment 1: OSEM_spectra.pdf  158 kB  | Hide | Hide all
OSEM_spectra.pdf
ELOG V3.1.3-